OpenStudy (gahm8684):

If the quantity 4 times x times y cubed plus 8 times x squared times y to the fifth power all over 2 times x times y squared is completely simplified to 2xayb + 4xcyd, where a, b, c, and d represent integer exponents, what is the value of a? _______ Numerical Answers Expected!

OpenStudy (gahm8684):

@nikato

OpenStudy (gahm8684):

\[\frac{ 4xy^3+8x^2y^5 }{ 2xy^2 }\]

OpenStudy (gahm8684):

@amistre64 @Compassionate @campbell_st @SolomonZelman

OpenStudy (amistre64):

what does: Numerical Answers Expected mean?

OpenStudy (amistre64):

\[\frac{ 4xy^3+8x^2y^5 }{ 2xy^2 }=2x^ay^b + 4x^cy^d\] just simplify and compare the parts

OpenStudy (gahm8684):

dont get it:c